OpenStudy (samigupta8):

if the equation x^2-cx+d-0 has roots equal to the fourth power of the roots of x^2+ax+b=0 where a^2>4b then the roots of x^2-4bx+2b^2-c=0 will be a.both real b.both negative c.both positive d.one positive and one negative

OpenStudy (samigupta8):

@hartnn pls..hlp

OpenStudy (raden):

let the roots of x^2-cx+d=0 is α and β while the roots of x^2+ax+b=0 is p and q. according information above, known α = p^4 and β = q^4. now from the equation of x^2-cx+d=0 we get α + β = c while from the equation of x^2+ax+b=0 we get p+q = c and pq = b because α = p^4 and β = q^4. so, α + β = p^4 + q^4. simplied α + β = (p^2 + q^2)^2 - 2(pq)^2 α + β = ((p + q)^2 - 2pq)^2 - 2(pq)^2 distribute the values of α + β = c, p+q=c and pq = b, we get α + β = ((-a)^2 - 2b)^2 - 2(b)^2 c = a^4 - 4a^2b + 4b^2 - 2b^2 c = a^4 - 4a^2b + 2b^2 now look at the equation x^2-4bx+2b^2-c=0, by complete square it can be ; x^2-4bx+2b^2-c=0 (x - 2b)^2 - 2b^2 - c = 0 (x - 2b)^2 = 2b^2 + c again distribute that c = a^4 - 4a^2b + 2b^2, therefore (x - 2b)^2 = 2b^2 + c (x - 2b)^2 = 2b^2 +a^4 - 4a^2b + 2b^2 (x - 2b)^2 = a^4 - 4a^2b + 4b^2 complete square again right side : (x - 2b)^2 = (a^2 - 2b)^2 finally, we get x : x - 2b = +- sqrt((a^2 - 2b)^2) x - 2b = +- (a^2 - 2b) x = 2b +- (a^2 - 2b) x1 = 2b + a^2 - 2b = a^2 x2 = 2b - a^2 + 2b = 4b - a^2 now, in logically. the value a^2 always be positive. while the value of 4b - a^2 would be negative. (because information above says a^2 > 4b)

OpenStudy (raden):

so, the answer is D. one positive and one negative

OpenStudy (anonymous):

\[ x^2-cx+d=0 \ \ ; \ \ (x-\alpha)(x-\beta) \\ \ \hspace{125 px} \alpha + \beta = c\] \[x^2+ax+b=0 \ \ ; \ \ (x+\sigma)(x+\lambda) \\ \hspace{125px} \sigma \lambda = b \\ \hspace{125px} \alpha = \sigma^4 \\ \hspace{125px} \beta = \lambda^4\] \[x^2-4bx+2b^2-c=0 \\ \hspace{125px} D = 16b^2 - 8b^2+4c \\ \hspace{125px} \ \ = 8b^2 + 4c \\ \hspace{125px} \ \ = 8 (\sigma \lambda)^2 + 4(\sigma^4 + \lambda^4)\] \[ \textrm{roots are} \ \ \frac{- \sigma \lambda ±\sqrt{8 (\sigma \lambda)^2 + 4(\sigma^4 + \lambda^4)}}{2}\] \[ D > -\sigma \lambda\] So yeah, what RaDen said. It's D Just wanted to finish typing it :P

OpenStudy (anonymous):

~~~ roots are \[2 \sigma \lambda ± \sqrt{2\sigma^2 \lambda^2 + \sigma^4 + \lambda^4}\]

OpenStudy (samigupta8):

i think that p+q shud be -a hw u wrote p+q=c

OpenStudy (raden):

sorry, that should p+q = -a not c. but the next step already be correct